OpenStudy (anonymous):

Give the exact value of the expression without using a calculator: sin(2cos(inv) 1/5)

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

let cos(inv) 1/5 = A

OpenStudy (anonymous):

\[\sin (2A)= 2sinAcosA\]

OpenStudy (anonymous):

what is A you substitute

OpenStudy (anonymous):

\[2\sin (\cos ^{-1}1/5) \cos (\cos ^{-1}1/5)\]

OpenStudy (anonymous):

\[2\sin (\cos ^{-1}1/5) * 1/5\]

OpenStudy (anonymous):

can you go on from here